OpenStudy (anonymous):

(i) Show that the equation 2 tan^2 θ sin^2 θ = 1 can be written in the form 2 sin^4 θ + sin^2 θ − 1 = 0.

OpenStudy (lgbasallote):

uhhh there's no over cos^2 on the bottom equation?

OpenStudy (anonymous):

Tan= Sin/Cos Try messing with that.

OpenStudy (ash2326):

We have \[2 \tan^2 \theta\ \sin^2 \theta=1\] We know that \[\tan \theta= \frac{\sin \theta}{\cos \theta}\] so \[2 \sin ^2 \theta \sin^2 \theta= \cos ^2 \theta\] Write \[\cos^2 \theta= 1- \sin^2 \theta\] I think you can do it from here:)

OpenStudy (anonymous):

@ash2326 Good job! So it the last part call the cos identity?

OpenStudy (anonymous):

Yes, I have it.

OpenStudy (ash2326):

@Romero Thanks but I don't know what's it's called it's a standard identity!!

OpenStudy (anonymous):

I was just asking because I always encounter the half angle identity and I thought his had a specific name to it.

OpenStudy (anonymous):

\[{2\sin^2 \theta \sin^2 \theta \over \cos^2 \theta}=1\]

OpenStudy (anonymous):

\[2\sin^4 \theta = 1-\cos^2 \theta\]

OpenStudy (anonymous):

\[2\sin^4 \theta = 1 -\sin^2 \theta\]

OpenStudy (anonymous):

which brings it back to it's origin

OpenStudy (anonymous):

I mean = cos^ theta. which gives 1- sin^2 theta

OpenStudy (lgbasallote):

uhmm why is it 1 - cos^2 theta? i think what you did was cross multiplication...i dunno

OpenStudy (anonymous):

I made a mistake there... I wrote cos^2 theta which gives 1-sin^2 instead again,. Hope that makes sense.

OpenStudy (lgbasallote):

oh i see good job :)